OpenStudy (anonymous):

Find the real number x satisfying \[\log_{1/2} (2x + 1 ) <-3\]?

OpenStudy (anonymous):

this type was done by @satellite73 yesterday but am confused

OpenStudy (anonymous):

@satellite73 can u help me out

OpenStudy (anonymous):

yes lets change the base to make life easier

OpenStudy (anonymous):

\[\log_{\frac{1}{2}}(x)=-\log_2(x)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if that is not clear, lets show it by the change of base formula we know \[\log_{\frac{1}{2}}(x)=\frac{\log_2(x)}{\log_2(\frac{1}{2})}=\frac{\log_2(x)}{-1}=-\log_2(x)\]

OpenStudy (anonymous):

site is stuck

OpenStudy (anonymous):

ok now we can solve \[-\log_2(2x+1)<-3\] easily change the signs and the sense of the inequality and get \[\log_2(2x+1)>3\]

OpenStudy (anonymous):

since log base 2 is an increasing function we conclude that \[2x+1>2^3\] \[2x+1>8\] and solve that one

OpenStudy (anonymous):

2x+1=8 x=7/2 is it....

OpenStudy (anonymous):

it is an inequality, not an equality

OpenStudy (anonymous):

\(x>\frac{7}{2}\)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

hope steps were clear

OpenStudy (anonymous):

hey wat is the difference between the yesterdays question and todyss can u tell plz

OpenStudy (anonymous):

\[ \log_{1/3} (\log_{4} (x^2-5)>0\]

OpenStudy (anonymous):

@satellite73